OpenStudy (anonymous):

Factor completely 12t^8 - 75t^4

OpenStudy (michele_laino):

plase put in evidence 3t^4, first

OpenStudy (michele_laino):

you should get this: \[3t ^{4}(4t ^{4}-25)=3t ^{4}(2t ^{2}-5)(2t ^{2}+5)=\] \[=3t ^{4}(\sqrt{2}t-\sqrt{5})(\sqrt{2}t+\sqrt{5})(2t ^{2}+5)\]

OpenStudy (anonymous):

I don't get the last oart

OpenStudy (anonymous):

Part *

OpenStudy (michele_laino):

the last part, is because: \[2t ^{2}-5\] is a difference between two square quantities, namely: \[\sqrt{2}t, \sqrt{5}\] so you have to apply the subsequent formula, as you well know: \[a ^{2}-b ^{2}=(a+b)(a-b)\] where: \[a=\sqrt{2}t, b=\sqrt{5}\]

OpenStudy (anonymous):

I never leArned that this is algebra 2 right?

OpenStudy (michele_laino):

I think so, even if you can stop at first step

OpenStudy (anonymous):

What about this ques. Express the product of (1/2y^2 -1/3y) and (12y+3/5) as a trinomial

OpenStudy (michele_laino):

is your product this? \[(\frac{ 1 }{ 2 }y ^{2}-\frac{ 1 }{ 3 }y)(12y+\frac{ 3 }{ 5 })\]

OpenStudy (anonymous):

Yes!

OpenStudy (michele_laino):

ok! you have to apply the distributive property of multiplication over the sum:

OpenStudy (michele_laino):

sorry I rewrite better: \[\frac{ y ^{2} }{ 2 }*\left( 12y+\frac{ 3 }{ 5 } \right)-\frac{ y }{ 3 }*\left( 12y+\frac{ 3 }{ 5 } \right)=\]

OpenStudy (michele_laino):

so, we have: \[\frac{ y ^{?} }{ 2 }*12y +\frac{ y ^{2} }{ 2 }*\frac{ 3 }{ 5 }-\frac{ y }{ 3 }*12y-\frac{ y }{ 3 }*\frac{ 3 }{ 5 } = \]

OpenStudy (michele_laino):

sorry the first term is: \[\frac{ y ^{2} }{ 2 }*12y\]

OpenStudy (michele_laino):

s after some simplification we get \[6y ^{3}-\frac{ 37 }{ 10 }y ^{2}-\frac{ y }{ 5 }\] @Sarah_s

OpenStudy (anonymous):

So that would be the answer?

OpenStudy (michele_laino):

yes, is it right?

OpenStudy (anonymous):

Okay and for this one If f(x)=x^2 -6, find f^-1(x) Is the answer y=x^2+6?

OpenStudy (michele_laino):

please do you want to calculate: \[\frac{ 1 }{ f(x) }\]

OpenStudy (anonymous):

No it's f^-1(x) so the inverse

OpenStudy (michele_laino):

ok, then inverse function is: \[x=\pm \sqrt{y+6}\]

OpenStudy (anonymous):

How

OpenStudy (michele_laino):

please insert x from the above formula into expression for f(x) and you will get y

OpenStudy (anonymous):

Wouldn't the inverse be x^2-6 then get in y = for! by +6?